OpenStudy (anonymous):

9^x+5 = 27^x-13 solve using Change of base

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

\[9^{x+5}=27^{x-13}\] Am i correct ?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Oh, good catch, the other problem is a bit uglier :-)

OpenStudy (anonymous):

\[\left( 3^{2} \right)^{x+5}=\left( 3^{3} \right)^{x-13}\] can you solve it ?

OpenStudy (whpalmer4):

using the requested change of base, rewrite the equation so that you have the same base on both sides

OpenStudy (anonymous):

you distrube?

OpenStudy (whpalmer4):

remember that \[(ab)^n = a^nb^n\]

OpenStudy (whpalmer4):

\[3^2=3*3\]\[3^3=3*3*3\]

OpenStudy (anonymous):

\[3^{2x+10}=3^{3x-39}\]

OpenStudy (anonymous):

x=49?

OpenStudy (whpalmer4):

Very good!

OpenStudy (whpalmer4):

the "other" equation actually has a solution of \(x=1\) but the other two are pretty ugly, and involve the logs of complex numbers :-)

OpenStudy (anonymous):

can you help me @whpalmer4 with that

OpenStudy (whpalmer4):

"other" equation being \[9^x+5 = 27^x-13\]

OpenStudy (anonymous):

can you help me with another probkem

OpenStudy (anonymous):

problem?

OpenStudy (whpalmer4):

I can try, what's the problem?

OpenStudy (anonymous):

solve for x 10^3a-1 = 10^7a+11

OpenStudy (whpalmer4):

there's no \(x\) there :-)

OpenStudy (whpalmer4):

\[10^{3a-1} = 10^{7a+11}\]Solve for \(a\) Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay, as you have two exponentials with the same base and nothing else, isn't it clear that \[3a-1=7a+11\]?

OpenStudy (whpalmer4):

We could take the \(\log_{10}\) of both sides to get that, if you wanted to be formal about it.

OpenStudy (anonymous):

ok